OpenStudy (sydney3545):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

jimthompson5910 (jim_thompson5910):

Start with f(x) = 9x - 2 and replace every x with g(x)

jimthompson5910 (jim_thompson5910):

so, f(x) = 9x - 2 f(g(x)) = 9(g(x)) - 2

jimthompson5910 (jim_thompson5910):

since we have f(g(x)) = 9(g(x)) - 2 and because g(x) = -x+3, we know that f(g(x)) = 9(g(x)) - 2 f(g(x)) = 9(-x+3 ) - 2... replace g(x) with -x+3 I'll let you simplify

OpenStudy (sydney3545):

im getting B @jim_thompson5910

jimthompson5910 (jim_thompson5910):

9(-x+3 ) - 2 = 9(-x)+9(3 ) - 2 9(-x+3 ) - 2 = -9x + 27 - 2 9(-x+3 ) - 2 = ???

OpenStudy (sydney3545):

Im lost what to do after this point i keep getting 6 but that now a choice

jimthompson5910 (jim_thompson5910):

do you see how I got -9x + 27 - 2

OpenStudy (sydney3545):

yes

jimthompson5910 (jim_thompson5910):

so you just need to simplify -9x + 27 - 2

jimthompson5910 (jim_thompson5910):

by combining like terms

OpenStudy (sydney3545):

so combine the -9x right?

jimthompson5910 (jim_thompson5910):

nope, the 27 and -2 are like terms

jimthompson5910 (jim_thompson5910):

the only like terms

OpenStudy (sydney3545):

So then it would be -9x+25

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (sydney3545):

can you help me with one more? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure, one more

OpenStudy (sydney3545):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

jimthompson5910 (jim_thompson5910):

same idea as last time

OpenStudy (sydney3545):

I tried doing the same thing as the last one and got -4x+-17

jimthompson5910 (jim_thompson5910):

f(x) = -4x + 7 f(g(x)) = -4(g(x)) + 7 ... replace every x with g(x) f(g(x)) = -4(10x - 6) + 7 ... replace the g(x)'s on the right side with 10x-6 f(g(x)) = -4(10x) - 4(-6) + 7 f(g(x)) = -40x + 24 + 7 \(\Large \leftarrow \text{see note below}\) f(g(x)) = -40x + 31 Note: this is where you made a mistake. Notice how the -4 outside multiplies with the -6 to get +24 and NOT -24.

OpenStudy (sydney3545):

Thanks !

jimthompson5910 (jim_thompson5910):

you're welcome